a) Sea water contains roughly 28.0 g of NaCl per liter. (NaCl molar mass = 58.44 gmol-1).

i. Calculate the number of moles of NaCl in a liter of sea water.

ii. Calculate the molarity of NaCl in sea water.

iii. Calculate the mass by volume percent (W/V) of NaCl in sea water.

b) To identify the cations and anions present in an unknown solution, the following tests were carried out.

I. A flame test on the solid substance produces a brick red colour.

II. Nitric acid was added to a solution of the substance followed by a few drops

of barium nitrate solution. A white precipitate was formed.

i. What is the cation present in the substance?

ii. What is the anion present in the substance?

iii. Give the formula of the substance present.


1
Expert's answer
2021-09-17T02:13:09-0400

Mass in a sea water = 28.0g per liter

Molar = 58.44

= 28.0/58.44 = 0.479moles

Molarity = moles/L

= 0.479/1

0.479 mol/L


Mass by volume percentage = 28/1000 × 100 = 2.8%